# Summary of the NOTSOFAR-1 Challenge: Highlights and Learnings

### v1

@@ Abstract
The first Natural Office Talkers in Settings of Far-field Audio Recordings (NOTSOFAR-1) Challenge is a pivotal initiative that sets new benchmarks by offering datasets more representative of the needs of real-world business applications than those previously available. The challenge provides a unique combination of 280 recorded meetings across 30 diverse environments, capturing real-world acoustic conditions and conversational dynamics, and a 1000-hour simulated training dataset, synthesized with enhanced authenticity for real-world generalization, incorporating 15,000 real acoustic transfer functions. In this paper, we provide an overview of the systems submitted to the challenge and analyze the top-performing approaches, hypothesizing the factors behind their success. Additionally, we highlight promising directions left unexplored by participants. By presenting key findings and actionable insights, this work aims to drive further innovation and progress in DASR research and applications.
@@ 1 Introduction
Advancements in the DASR field have been long supported by comprehensive and high-quality datasets. For example, the AMI dataset [1] provides 100 hours of real English meeting recordings with close-microphone annotations, which has been widely used as a benchmark for DASR. More recently, the CHiME-5/6 challenge [2] [3] provided about 50 hours of daily conversation recordings along with baseline DASR systems, significantly advancing DASR research. The LibriCSS dataset [4] provides evaluation data on more controlled environments by playing LibriSpeech [5] utterances via loudspeakers. Nevertheless, there are still significant gaps between real-world complexity, such as limited recording environments (AMI), fixed numbers of speakers (AMI, CHiME, LibriCSS), etc. These gaps pose challenges to developing a DASR system that works under real-world uncertainty.
To address this, the Natural Office Talkers in Settings of Far-field Audio Recordings (NOTSOFAR-1) challenge [6], an international competition in DASR, was introduced at Interspeech 2024 as part of CHiME’s annual challenge (CHiME-8 [7]). It provided comprehensive datasets as well as a baseline system to facilitate research and development in the field of DASR. The primary goals of the NOTSOFAR-1 challenge were to advance the state-of-the-art in DASR by offering new datasets that capture the complexities of distant meeting transcription and to provide a platform for researchers to evaluate their systems in realistic scenarios. The challenge aimed to bridge the gap between existing datasets and real-world applications, focusing on single-channel and known-geometry multi-channel tracks, which are commonly used in conference room environments. It introduced two new datasets: a benchmark dataset of 280 English meetings recorded across 30 different rooms, and a 1000-hour simulated training dataset carefully designed to generalize well to real-world scenarios. The simulated dataset incorporates 15,000 real acoustic transfer functions (ATFs) [8] and features high-quality clean speech, providing a rich resource for training speech separation and enhancement methods.
@@ 2.2 Datasets
The challenge introduced two new datasets to the community:
Recorded Meeting Dataset: Comprises 280 unique English meetings with high-quality transcription, each of which averages six minutes in duration, recorded in 30 different rooms with 4-8 attendees per meeting. The recordings capture a broad spectrum of acoustic and conversational patterns, providing a rich resource for benchmarking DASR systems. The data set includes detailed metadata (e.g., moving speaker, crowded meeting, etc.) for error analysis, and employs both SC and MC devices.
@@ 2.3 Metrics
Time-constrained minimum-permutation word-error rate (tcpWER): Measures the word error rate while attributing errors to specific speakers within a time-constrained window. It provides an assessment of both speech recognition and speaker diarization performance.
Time-constrained optimal reference combination word-error rate (tcorcWER): Optimizes word error rate calculation by dynamically selecting the most accurate reference transcriptions within specific time windows. It focuses on the accuracy of the transcribed text, regardless of speaker identity.
@@ 2.5 Results
We received 9 submissions in the single-channel track and 5 NOTSOFAR-1 submissions in the multi-channel track. We’ve also included 3 DASR track submissions to the multi-channel track (see 2.4 for more details). Most of the submissions improved the baselines, some very significantly. Tables 1 and 2 show tcpWER and tcorcWER for all submissions to the SC and MC tracks respectively.
@@ 3 Overview of the submitted systems
  • CSS-ASR-Dia pipelines start with continuous source separation (CSS) [4] as a front-end, followed by automatic speech recognition (ASR) and speaker diarization (Dia). CSS module generates a set of non-overlapped speech signals from a continuous audio stream that contains multiple utterances that are partially overlapped by a varying degree. Speech signals produced by CSS are processed by ASR. Since the speech signals generated by CSS can combine multiple speakers (the only guarantee is that none of the signals includes overlapped speech), ASR is followed by speaker diarization.
@@
As challenge designers interested in encouraging participants to explore the opportunities of replacing traditional GSS-based approaches to source separation with data-driven neural-network based alternatives, a CSS-ASR-Dia pipeline was provided as a baseline system. We also wanted to give participants the option to adapt the ASR to the distant speech as opposed to improving the acoustic front-end in order to bridge the gap between distant and close, so the ASR module was not fixed and participants were encouraged to come up with alternatives.
@@
Tables 3 and 4 show the techniques used by MC and SC systems respectively. Participants could submit up to four systems, and the tables include all reported techniques, not just those from the best-performing solutions. Submissions that underperformed compared to the baseline on the evaluation set are excluded.
@@ 4.2 Diarization and ASR
GSS depends on accurate speaker diarization, so all three teams developed sophisticated diarization subsystems combining clustering-based speaker diarization (CSD), neural speaker diarization (NSD), and speech separation components to refine speaker embedding in regions with overlapping speech. All leading diarization systems incorporated TS-VAD [19] techniques, specifically leveraging NSD-MS2S [20]. Given the importance of precise speaker counting for TS-VAD initialization, teams implemented advanced speaker-counting modules. STCON used a combination of an in-house version of Wav2vec 2.0 [10] [21] trained to extract several embeddings for segments with mixed speech and SpeechBrain ECAPA-TDNN model [22]. NTT speaker counting module employed ECAPA-TDNN speaker embeddings derived from on the audio processed by GSS initialized with EEND-based segmentation [23] [11]. USTC system ran CSS on the overlapping speech segments before calculating speaker embeddings [12].
@@
ASR modules of the top system also received substantial upgrades. USTC introduced “Enhanced Whisper” leveraging features extracted from self-supervised pre-trained models like WavLM [24]. They also introduced additional improvements like augmentation of the original Whisper model [25] with a ConvNeXt structure [26], running in parallel to the standard 1D convolutions, adopted bias relative positional encoding, integrated sigmoid gating mechanism augmented the final layer of the encoder with a Mixture of Experts (MoE) component. STCON used an ensemble of DNN-HMM hybrid models and end-to-end (E2E) models trained on GSS-processed CHiME-8 train datasets and the clean part of LibriSpeech [5] distorted with the simulated room impluse responses (RIR) selected to match those detected in the CHiME-8 training data. NTT used an ensemble of fine-tuned Whisper Large v3, Whisper Medium, NeMo Transducer [27] and fine tuned WavLM Transducer [24] [28]. The team also built a Transformer-LM used to rescore the n-best hypotheses of each model and combined the results using ROVER [29].
@@
With all top teams investing heavily into multiple components, it is interesting to explore how the accuracy of each component affects the the final speech recognition accuracy. While drawing decisive conclusions is impossible without thorough hands-on experimentation with the systems, the similar structure of the top systems and the ablation studies provided by the teams offer valuable insights into contribution of specific components into USTC’s system success.
@@
Since all top systems relied on GSS for source separation and the improvements to the separation component proposed by the winning team (GSS initialization in T-F domain based on mask prediction by JDS [12] didn’t result in a major improvement based on the numbers reported by the team (Table 2 in [12]), we take the risk of assuming no significant difference in the systems source separation components and exclude the source separation from the discussion to focus on the accuracy of diarization subsystems.
@@ 4.3 Spatial information for diarization
Although every submitted multi-channel system used that spatial info for speech separation through GSS [9] or Minimum Variance Distorshionless Response (MVDR) beamformer [31], most of the systems ignored it for the diarization task and processed the audio channels independently benefiting only from aggregating multiple versions of diarization.
@@
We believe the existence of well-established neural net based approaches to speaker diarization based on processing of a single-channel audio (EEND-VC [23], TSVAD [19], NSD-MS2S [20]) and the lack of such well established approaches for multi-channel audio is the reason. We hope that better use of the spatial signal offered by microphone arrays still presents an opportunity for a further significant improvement in multi-channel speaker diarization.
@@ 4.4 Beyond GSS
Table 2 of USTC paper [12] shows tcpWER for every combination of their diarization and separation subsystems. Columns correspond to diarization systems (Sys-1 RTTM to Sys-4 RTTM), rows - to separation systems (V1 to V3). V1 skips JDS initializing GSS in time domain, V3 skips GSS with MVDR beamformer [31] calculated based on T-F masks predicted by JDS, while V2 uses JDS to initialize GSS in time-frequency domain. One can see that V2 performance is consistently but negligibly better than that of V1 showing that no significant accuracy improvement was achieved by T-F domain initialization of GSS. Besides accuracy improvement, T-F domain initialization could potentially reduce the number of iterations of GSS optimization algorithm required to achieve the desired quality of separation. Unfortunately, USTC team didn’t report the number of iterations they used for either of their system, so no such comparison is possible.
@@
V3 separation system by USTC replaces GSS-based source separation with MVDR beamformer [31] calculated based on by per speaker T-F masks predicted by a neural JDS module [12]. Comparing the performance of V1, V2 and V3 systems in Table 2 of the USTC report [12] provides insight into the ability of neural T-F mask prediction replace GSS. V3 is consistently worse than both V2 (including both JDS and GSS) and V1 (GSS only), with the relative degradation of 3 to 4% depending on the diarization. USTC team’s decision to use the predicted T-F masks to guide a beamformer rather than use them directly to reconstruct the source audio suggests the ASR accuracy on directly reconstructed masks was inferior. If this assumption is true, STCON might observe a similar phenomenon of higher recognition error rate on neurally separated speech in spite of high quality of the restored audio [10]. STCON’s discussion on Target Speaker Extractor (TSE) can provide some insight into the phenomenon. We provide the summary of that discussion in the following paragraph.
@@
Another potential approach to improvement of the classical GSS+beamforming source separation is replacing a linear beamformer with a neural Target Speaker Extraction (TSE) model. This approach was explored by STCON team [10]. They used the conformer architecture offered by the baseline CSS model, but extended it to process target speaker info together with mixture. According to the team’s report, standalone training of the TSE model didn’t yield better ASR accuracy than GSS+beamforming pipeline despite the good quality of target speaker’s speech and low SNR in most of the TSE outputs. The team attributed the gap in ASR accuracy to two factors: a mismatch between loss functions in TSE (si-sdr) and ASR (CTC+attention) and the longer context used by GSS compared to TSE. The first factor was addressed by concatenating TSE and ASR models and fine-tuning the pre-trained TSE model using the ASR loss. The second factor was tackled by utilizing the GSS output as a reference channel in the TSE output, effectively turning the TSE into a “GSS signal enhancer” which the team referred to as the Guided Target Speaker Extractor (G-TSE). These two changes helped bridge the gap and achieve improvements over the GSS+beamforming pipeline, supporting the team’s hypothesis.
@@ 4.5 CSS-based systems
CSS-based systems like those submitted by NPU-TEA [13] and NAIST [15] teams improved the baseline system by a lot, but failed to achieve the same transcription accuracy as Dia-Sep-ASR systems. The best tcpWER on the evaluation dataset being 18.7% (NPU-TEA) while the winning system achieved tcpWER=10.8% (USTC). The baseline system tcpWER is 28.3%.
Both NPU-TEA and NAIST used conformer architecture proposed by the baseline. The NAIST team reused the baseline CSS model, but added a de-reverberation module. Despite the CSS model being trained to predict reverberation-free signals, the NAIST team discovered reverberation in the model output. Their experiments shown that the added a Weighted Prediction Error (WPE) module was responsible for reducing tcpWER from 30.3% to 28.9% on dev-set-2 (Figure 4 in [15].)
@@
Although teams extensively used the provided real training data for fine-tuning their diarization and ASR model, the creators of the CSS-based systems relied solely on training with simulated data, leaving fine-tuning of CSS with real meetings data unexplored. Although, as the challenge designers, we invested significant efforts into realistic simulation of far-field audio and natural meeting dynamic, we believe the remaining gap between synthetic and real data may have prevented CSS from reaching its full potential. We encourage further exploration of CSS fine-tuning with real data and hope that bridging the data gap will unlock to the full capabilities of neural network-based CSS.

### v2

@@ Abstract
The first Natural Office Talkers in Settings of Far-field Audio Recordings (NOTSOFAR-1) Challenge is a pivotal initiative that sets new benchmarks by offering datasets more representative of the needs of real-world business applications than those previously available. The challenge provides a unique combination of 315 recorded meetings across 30 diverse environments, capturing real-world acoustic conditions and conversational dynamics, and a 1000-hour simulated training dataset, synthesized with enhanced authenticity for real-world generalization, incorporating 15,000 real acoustic transfer functions. In this paper, we provide an overview of the systems submitted to the challenge and analyze the top-performing approaches, hypothesizing the factors behind their success. Additionally, we highlight promising directions left unexplored by participants. By presenting key findings and actionable insights, this work aims to drive further innovation and progress in DASR research and applications.
@@ 1 Introduction
Advancements in the DASR field have been long supported by comprehensive and high-quality datasets. For example, the AMI dataset [1] provides 100 hours of real English meeting recordings with close-microphone annotations, which has been widely used as a benchmark for DASR. More recently, the CHiME-5/6 challenge [2, 3] provided about 50 hours of daily conversation recordings along with baseline DASR systems, significantly advancing DASR research. The LibriCSS dataset [4] provides evaluation data on more controlled environments by playing LibriSpeech [5] utterances via loudspeakers. Nevertheless, there are still significant gaps between real-world complexity, such as limited recording environments (AMI), fixed numbers of speakers (AMI, CHiME, LibriCSS), etc. These gaps pose challenges to developing a DASR system that works under real-world uncertainty.
To address this, the Natural Office Talkers in Settings of Far-field Audio Recordings (NOTSOFAR-1) challenge [6], an international competition in DASR, was introduced at Interspeech 2024 as part of CHiME’s annual challenge (CHiME-8 [7])  [^3] A more detailed comparison between NOTSOFAR-1 and other benchmarks can be found in [6, Section 1]. It provided comprehensive datasets as well as a baseline system to facilitate research and development in the field of DASR. The primary goals of the NOTSOFAR-1 challenge were to advance the state-of-the-art in DASR by offering new datasets that capture the complexities of distant meeting transcription and to provide a platform for researchers to evaluate their systems in realistic scenarios. The challenge aimed to bridge the gap between existing datasets and real-world applications, focusing on single-channel and known-geometry multi-channel tracks, which are commonly used in conference room environments. It introduced two new datasets: a benchmark dataset of 315 English meetings recorded across 30 different rooms, and a 1000-hour simulated training dataset carefully designed to generalize well to real-world scenarios. The simulated dataset incorporates 15,000 real acoustic transfer functions (ATFs) [8] and features high-quality clean speech, providing a rich resource for training speech separation and enhancement methods.
@@ 2.2 Datasets
The challenge introduced two new datasets to the community–a recorded meeting dataset and a simulated training dataset. We describe the key attributes of each dataset below and refer the reader to [6, Section 4] for comprehensive details. [^4] Further details of the datasets, including metadata analysis, are available on the challenge’s website: https://www.chimechallenge.org/challenges/chime8/task2/data.
Recorded Meeting Dataset: Comprises 315 unique English meetings with high-quality transcription, each of which averages six minutes in duration, recorded in 30 different rooms with both SC and MC devices. Each meeting features 4-8 adult participants, with 22 unique speakers participated in the development and training sets, and a separate set of 13 speakers in the evaluation set.
The transcription process included a multi-judge annotation, and avoided machine assistance to minimize potential biases. The recordings capture a broad spectrum of acoustic and conversational patterns, e.g., different topics, moving speakers, un/crowded meetings, overlap speech, etc., where part of this information is provided as metadata for error analysis.
For completeness, Table 1 presents the summary table from [6, Table 1], which outlines the technical details of this dataset.
Feature Training Set Development Set Evaluation Set (blind)
Number of Meetings 110 meetings 35 meetings 170 meetings
Average Meeting Duration 6 minutes 6 minutes 6 minutes
Number of SC streams 5 streams 5 streams 6 streams
Number of MC devices 4 devices 4 devices 4 devices
Total Duration of SC Audio Recordings 55 hours (110 meetings $\times$× 6 minutes $\times$× 5 SC streams) 17.5 hours (35 meetings $\times$× 6 minutes $\times$× 5 SC streams) 102 hours (170 meetings $\times$× 6 minutes $\times$× 6 SC streams)
Total Duration of MC Audio Recordings 44 hours (110 meetings $\times$× 6 minutes $\times$× 4 MC devices) 14 hours (35 meetings $\times$× 6 minutes $\times$× 4 MC devices) 68 hours (170 meetings $\times$× 6 minutes $\times$× 4 MC devices)
Total Number of Rooms 20 rooms 5 rooms 13 rooms
Number of Speakers 22 speakers 11 speakers 13 speakers
Language English
Table 1: A summarization of NOTSOFAR recorded meeting dataset as presented in [6, Table 1]. SC and MC stand for Single-Channel and Multi-Channel, respectively.
@@ 2.3 Metrics
Time-constrained minimum-permutation word-error rate (tcpWER) [10]: A time-constrained version of the Concatenated minimum-Permutation WER (cpWER) [3]. It measures word error rate while attributing errors to specific speakers within a time-constrained window, providing an assessment of both speech recognition and speaker diarization performance.
Time-constrained optimal reference combination word-error rate (tcorcWER) [10]: A time-constrained version of the Optimal Reference Combination WER (orcWER) [11]. It optimizes word error rate calculation by dynamically selecting the most accurate reference transcriptions within specific time windows, focusing solely on the accuracy of the transcribed text without considering speaker identity.
@@ 2.5 Results
We received 9 submissions in the single-channel track and 5 NOTSOFAR-1 submissions in the multi-channel track. We’ve also included 3 DASR track submissions to the multi-channel track (see 2.4 for more details). Most of the submissions improved the baselines, some very significantly. Tables 2 and 3 show tcpWER and tcorcWER for all submissions to the SC and MC tracks respectively.
@@ 3 Overview of the submitted systems
  • CSS-ASR-Dia pipelines start with continuous source separation (CSS) [4] as a front-end, followed by automatic speech recognition (ASR) and speaker diarization (Dia). CSS module generates a set of non-overlapped speech signals from a continuous audio stream that contains multiple utterances that are partially overlapped by a varying degree. Speech signals produced by CSS are processed by ASR. Since the speech signals generated by CSS can combine multiple speakers – CSS aims to “unmix” overlapping speech rather than isolate individual speakers – ASR is followed by speaker diarization.
@@
As challenge designers, we aimed to encourage participants to explore replacing traditional GSS-based approaches to source separation with data-driven neural-network alternatives. This led us to provide a CSS-ASR-Dia pipeline as the baseline system. We also wanted to give participants the option to adapt the ASR to the distant speech as opposed to improving the acoustic front-end in order to bridge the gap between distant and close, so the ASR module was not fixed and participants were encouraged to come up with alternatives.
@@
Tables 4 and 5 show the techniques used by MC and SC systems respectively. Participants could submit up to four systems, and the tables include all reported techniques, not just those from the best-performing solutions. Submissions that underperformed compared to the baseline on the evaluation set are excluded.
@@ 4.2 Diarization and ASR
GSS depends on accurate speaker diarization, so all three teams developed sophisticated diarization subsystems combining clustering-based speaker diarization (CSD), neural speaker diarization (NSD), and speech separation components to refine speaker embedding in regions with overlapping speech. All leading diarization systems incorporated TS-VAD [21] techniques, specifically leveraging NSD-MS2S [22]. Given the importance of precise speaker counting for TS-VAD initialization, teams implemented advanced speaker-counting modules. STCON used a combination of an in-house version of Wav2vec 2.0 [12, 23] trained to extract several embeddings for segments with mixed speech and SpeechBrain ECAPA-TDNN model [24]. NTT speaker counting module employed ECAPA-TDNN speaker embeddings derived from on the audio processed by GSS initialized with EEND-based segmentation [25, 13]. USTC system ran CSS on the overlapping speech segments before calculating speaker embeddings [14].
@@
ASR modules of the top system also received substantial upgrades. USTC introduced “Enhanced Whisper” leveraging features extracted from self-supervised pre-trained models like WavLM [26]. They also introduced additional improvements like augmentation of the original Whisper model [27] with a ConvNeXt structure [28], running in parallel to the standard 1D convolutions, adopted bias relative positional encoding, integrated sigmoid gating mechanism augmented the final layer of the encoder with a Mixture of Experts (MoE) component. STCON used an ensemble of DNN-HMM hybrid models and end-to-end (E2E) models trained on GSS-processed CHiME-8 train datasets and the clean part of LibriSpeech [5] distorted with the simulated room impluse responses (RIR) selected to match those detected in the CHiME-8 training data. NTT used an ensemble of fine-tuned Whisper Large v3, Whisper Medium, NeMo Transducer [29] and fine tuned WavLM Transducer [26, 30]. The team also built a Transformer-LM used to rescore the n-best hypotheses of each model and combined the results using ROVER [31].
@@
With all top teams investing heavily into multiple components, it is interesting to explore how the accuracy of each component affects the final speech recognition accuracy. While drawing decisive conclusions is impossible without thorough hands-on experimentation with the systems, the similar structure of the top systems and the ablation studies provided by the teams offer valuable insights into contribution of specific components into USTC’s system success.
@@
Since all top systems relied on GSS for source separation and the improvements to the separation component proposed by the winning team (GSS initialization in T-F domain based on mask prediction by JDS [14]) didn’t result in a major improvement based on the numbers reported by the team (Table 2 in [14]), we take the risk of assuming no significant difference in the systems source separation components and exclude the source separation from the discussion to focus on the accuracy of diarization subsystems.
@@ 4.3 Spatial information for diarization
Although every submitted multi-channel system used that spatial info for speech separation through GSS [9] or Minimum Variance Distortionless Response (MVDR) beamformer [33], most of the systems ignored it for the diarization task and processed the audio channels independently benefiting only from aggregating multiple versions of diarization.
@@
We believe the existence of well-established neural net based approaches to speaker diarization based on processing of a single-channel audio (EEND-VC [25], TSVAD [21], NSD-MS2S [22]) together with the lack of such well established approaches for multi-channel audio is the reason for the dominance of single-channel oriented diarization subsystems in the challenge submissions. We hope that better use of the spatial signal offered by microphone arrays still presents an opportunity for a further significant improvement in multi-channel speaker diarization.
@@ 4.4 Beyond GSS
Table 2 of USTC paper [14] presents tcpWER for different combinations of diarization and separation subsystems. The results show that the separation subsystem with T-F domain initialization of GSS (V2) performed only marginally better than the same subsystem without T-F domain initialization (V1) across all diarization front-ends. This suggests that T-F domain initialization of GSS doesn’t significantly improve accuracy. However, beyond accuracy, T-F domain initialization may reduce the number of iterations required for GSS optimization to achieve the desired separation quality. Unfortunately, the USTC team didn’t report the number of iterations used ini their systems, making such comparison impossible.
@@
V3 separation system by USTC replaces GSS-based source separation with MVDR beamformer [33] calculated based on by per speaker T-F masks predicted by a neural JDS module [14]. Comparing the performance of V1, V2 and V3 systems in Table 2 of the USTC report [14] provides insight into the ability of neural T-F mask prediction to replace GSS. V3 is consistently worse than both V2 (including both JDS and GSS) and V1 (GSS only), with the relative degradation of 3 to 4% depending on the diarization. USTC team’s decision to use the predicted T-F masks to guide a beamformer rather than use them directly to reconstruct the source audio suggests the ASR accuracy on directly reconstructed masks was inferior. If this assumption is true, STCON might observe a similar phenomenon of higher recognition error rate on neurally separated speech in spite of high quality of the restored audio [12]. STCON’s discussion on Target Speaker Extractor (TSE) can provide some insight into the phenomenon. We provide the summary of that discussion in the following paragraph.
@@
Another potential approach to improvement of the classical GSS+beamforming source separation is replacing a linear beamformer with a neural Target Speaker Extraction (TSE) model. This approach was explored by STCON team [12]. They used the conformer architecture offered by the baseline CSS model, but extended it to process target speaker info together with mixture. According to the team’s report, standalone training of the TSE model didn’t yield better ASR accuracy than GSS+beamforming pipeline despite the good quality of target speaker’s speech and high SNR in most of the TSE outputs. The team attributed the gap in ASR accuracy to two factors: a mismatch between loss functions in TSE (si-sdr) and ASR (CTC+attention) and the longer context used by GSS compared to TSE. The first factor was addressed by concatenating TSE and ASR models and fine-tuning the pre-trained TSE model using the ASR loss. The second factor was tackled by utilizing the GSS output as a reference channel in the TSE output, effectively turning the TSE into a “GSS signal enhancer” which the team referred to as the Guided Target Speaker Extractor (G-TSE). These two changes helped bridge the gap and achieve improvements over the GSS+beamforming pipeline, supporting the team’s hypothesis.
@@ 4.5 CSS-based systems
CSS-based systems like those submitted by NPU-TEA [15] and NAIST [17] teams improved the baseline system by a lot, but failed to achieve the same transcription accuracy as Dia-Sep-ASR systems. The best tcpWER on the evaluation dataset achieved by a Dia-Sep-ASR system was 18.7% (NPU-TEA) while the winning system achieved tcpWER=10.8% (USTC). The baseline system tcpWER is 28.3%.
Both NPU-TEA and NAIST used conformer architecture proposed by the baseline. The NAIST team reused the baseline CSS model, but added a de-reverberation module. Despite the CSS model being trained to predict reverberation-free signals, the NAIST team discovered reverberation in the model output. Their experiments shown that the added Weighted Prediction Error (WPE) module was responsible for reducing tcpWER from 30.3% to 28.9% on dev-set-2 (Figure 4 in [17].)
@@
Although teams extensively used the provided real training data for fine-tuning their diarization and ASR model, the creators of the CSS-based systems relied solely on training with simulated data, leaving fine-tuning of CSS with real meetings data unexplored. Although, as the challenge designers, we invested significant efforts into realistic simulation of far-field audio and natural meeting dynamic, we believe the remaining gap between synthetic and real data may have prevented CSS from reaching its full potential. We encourage further exploration of CSS fine-tuning with real data and hope that bridging the data gap will unlock to the full capabilities of neural network-based CSS. We discuss this and other factors that could affect the performance of CSS-based system in the challenge in 6.